# ABACUS: An Electronic Structure Analysis Package for the AI Era

### v1

@@ III.1.3 Collinear and Non-Collinear Spin
$m(\mathbf{r})=\rho_{\uparrow}(\mathbf{r})-\rho_{\downarrow}(\mathbf{r}).$ (7)
The spin-dependent Kohn-Sham equation has effective potentials from the external, Hartree, and exchange-correlation potentials, with the latter now explicitly dependent on both the spin-up and spin-down densities.
The exchange-correlation potential $V_{\text{XC}}$ is split into two components $V_{\text{XC},\uparrow}(\mathbf{r})$ and $V_{\text{XC},\downarrow}(\mathbf{r})$, which are derived from the derivative of the exchange-correlation energy $E_{\text{XC}}[\rho_{\uparrow},\rho_{\downarrow}]$ with respect to the electron density
$V_{\text{XC},\sigma}(\mathbf{r})=\frac{\delta E_{\text{XC}}[\rho_{\uparrow},\rho_{\downarrow}]}{\delta\rho_{\sigma}(\mathbf{r})},$ (8)
where $\sigma$ represents either spin-up or spin-down electrons. Similarly to spinless cases, the approximations of $E_{\text{XC}}$ include the local spin density approximation (LSDA) Miehlich et al. (1989), where $E_{\text{XC}}$ depends only on the local spin densities. In addition, the generalized gradient approximation $E_{\text{XC}}$ also depends on the gradients of the spin densities. More sophisticated methods incorporate additional ingredients, such as the kinetic energy density for meta-GGA and occupied KS orbitals for hybrid functionals.
@@
The non-collinear spin calculation permits electron spins to orient in arbitrary directions rather than being restricted to a single axis. In the non-collinear spin KS-DFT, the Kohn-Sham electronic wave function $i$ can be represented as two-component spinors as
$\Psi_{i}(\mathbf{r})=\begin{pmatrix}\psi_{i\uparrow}(\mathbf{r})\\ \psi_{i\downarrow}(\mathbf{r})\end{pmatrix}.$ (9)
[…]
@@
$\left(\begin{array}[]{cc}\rho^{\uparrow\uparrow}&\rho^{\uparrow\downarrow}\\ \rho^{\downarrow\uparrow}&\rho^{\downarrow\downarrow}\end{array}\right)=(\rho\bm{I}+\bm{\sigma}\cdot\bm{m})=\frac{1}{2}\left(\begin{array}[]{cc}\rho+m_{z}&m_{x}-im_{y}\\ m_{x}+im_{y}&\rho-m_{z}\end{array}\right).$ (10)
@@
where $\bm{I}$ is $2*2$ diagonal matrix, $m_{i}|_{i=x,y,z}$ refer to magnetic density along three axis.
The exchange-correlation potential $V_{\rm xc}$ in the presence of non-collinear spins becomes a $2\times 2$ matrix that acts on these spinors
@@
where the off-diagonal terms $V_{\rm xc}^{\uparrow\downarrow}$ and $V_{\rm xc}^{\downarrow\uparrow}$ represent the non-collinear components of the potential, which mix the spin-up and spin-down states due to spin-orbit coupling (SOC) or other magnetic interactions.
@@
Taking advantage of the Pauli matrices, the exchange-correlation potential can be converted into:
$v_{xc}(\mathbf{r})=\frac{\delta E_{xc}}{\delta\rho(\mathbf{r})}=\frac{1}{2}(V_{xc}^{\uparrow\uparrow}(\mathbf{r})+V_{xc}^{\downarrow\downarrow}(\mathbf{r})),$ (12)
$b^{x}_{xc}(\mathbf{r})=\frac{\delta E_{xc}}{\delta m_{x}(\mathbf{r})}=\frac{1}{2}(V_{xc}^{\uparrow\downarrow}(\mathbf{r})+V_{xc}^{\downarrow\uparrow}(\mathbf{r})),$
$b^{y}_{xc}(\mathbf{r})=\frac{\delta E_{xc}}{\delta m_{y}(\mathbf{r})}=\frac{1}{2}(i(V_{xc}^{\uparrow\downarrow}(\mathbf{r})-V_{xc}^{\downarrow\uparrow})(\mathbf{r})),$
$b^{z}_{xc}(\mathbf{r})=\frac{\delta E_{xc}}{\delta m_{z}(\mathbf{r})}=\frac{1}{2}(V_{xc}^{\uparrow\uparrow}(\mathbf{r})-V_{xc}^{\downarrow\downarrow}(\mathbf{r})).$
Following this format, the charge density and exchange-correlation potential can be stored as real numbers rather than as complex numbers in the spinor format. When constructing the Hamiltonian, however, it is necessary to convert the exchange-correlation potential back into the spinor format to act on the two-component spinor wavefunctions. This approach facilitates the handling of non-collinear spin configurations within the DFT framework.
In ABACUS, the spin-orbit coupling effect are introduced through the non-local terms of the norm-conserving pseudopotentials using the Kleinman-Bylander (KB) projector Kleinman (1980); Bachelet and Schlüter (1982) formalism. The projector orbitals are expanded using complex spherical harmonics $Y_{lj}$ with orbital angular moment $l$ and total angular moment $j$ rather than the real spherical harmonics $Y_{lm}$. Given that the complex spherical harmonics $Y_{lj}$ can be expressed as combinations of the real spherical harmonics $Y_{lm}$, the formula of the non-local terms Hemstreet, Fong, and Nelson (1993); Theurich and Hill (2001) can be reformulated accordingly
$V_{NL}^{\sigma\sigma^{\prime}}=$ $\sum_{\tau}\sum_{pp^{\prime}}D_{\tau lj}\left[\sum_{m_{j}=-j}^{j}\alpha_{m_{j}}^{\sigma lj}U_{m_{j}m}^{\sigma lj}\alpha_{m_{j}}^{\sigma^{\prime}lj}U_{m_{j}m}^{*\sigma^{\prime}lj}\right]$ (13)
$\left|\beta_{\tau p}Y_{lm}\rangle\langle Y_{lm^{\prime}}\beta_{\tau p^{\prime}}\right|,$
where $U_{jm}^{\sigma lj}$ are unitary matrix and $\alpha_{m_{j}}^{\sigma lj}$ are Clebsch-Gordan coefficients, $D_{\tau lj}$ are radial pseudopotential components of non-local part with atom $\tau$ and $j=l+\frac{1}{2}$ for $l>=0$ and $j=l-\frac{1}{2}$ for $l>0$. Norm-Conserving pseudopotential with SOC effect is labeled as "full-relativistic" (FR), to distinguish them from "scalar-relativistic" (SR), which considers the relativistic mass correction without spin-orbit interactions.
III.2 Self-Consistent Field Method
@@
Figure 3: ABACUS performs self-consistent field calculations using two types of basis sets: PW and NAOs. The PW basis set employs iterative diagonalization methods like CG and Davidson to solve the Kohn-Sham equations. For NAOs, it uses exact diagonalization or the low-scaling method PEXSI. ABACUS applies different mixing algorithms for various types of calculations, including non-magnetic, collinear, and non-collinear. Additionally, ABACUS handles density matrix mixing for DFT+U, hybrid functionals, and DeePKS calculations with the NAOs basis.
The self-consistent field (SCF) method is an iterative method to update the electron density and solve the KS equation. For each electronic step during the SCF calculations, an input electron density $\rho_{\rm in}$ is used to construct the Hamiltonian, and an output electron density $\rho_{\rm out}$ is constructed from the Kohn-Sham wave functions. Typically, the self-consistency is achieved when the electron density is converged within a threshold.
Essentially, in order to obtain stable and efficient convergence during the SCF calculations for the nonlinear Kohn-Sham equation, several charge mixing methods have been proposed Pulay (1980, 1982); Johnson (1988). In this regard, a residual vector can be defined as
@@
The simplest mixing method is so-called plain mixing, where the input density of $(i+1)$-th iteration is determined by
$\rho^{i+1}_{in}=\rho^{i}_{in}+\alpha_{\rho}R[\rho^{i}_{in}],$ (15)
where $\alpha_{\rho}\in[0,1]$ specifies the linear mixing parameter.
Effective mixing algorithms aid the SCF procedure in correctly identifying the ground state, preventing entrapment in local minima, and avoiding convergence to excited states. Fig. 3 demonstrates that ABACUS incorporates sophisticated preconditioning and mixing strategies to guarantee stable and dependable SCF convergence. We will now provide a detailed discussion of these algorithms.
III.2.1 Modified Kerker Preconditioner
During SCF procedure, it is reasonable to assume that when the input charge density approaches the ground state charge density, the residual density relates proportionally to the discrepancy between the input charge density and the ground state charge density. In the context of linear response theory, the residual density is intricately associated with the dielectric function. The pace of convergence is primarily contingent upon $\overline{\Gamma}$, which signifies the energy spectrum of the dielectric function $\epsilon$ Kresse and Furthmüller (1996).
Mixing methods that forgo a preconditioner adopt a rudimentary dielectric matrix characterized by unity across its diagonal elements, that is, $\epsilon(q)=1$. This simplification may suffice for insulators or semiconductors where the dielectric function demonstrates gradual variation, in contrast to metals where it is subject to notable fluctuations. To address the issues with metals’ dielectric functions, a preconditioning approach, known as the Kerker preconditioner, incorporates a factor of $\frac{q^{2}}{q^{2}+q^{2}_{0}}$ for varying wavevector $\bm{q}$ to modulate the influence of distinct components Kresse and Furthmüller (1996), where $q=|\bm{q}|$ is the module of wavevector and $q_{0}$ is the damping parameter. Although this adjustment originates from the uniform electron gas model, it can significantly reduce long-range charge oscillations and accelerate convergence.
[…]
Excessive damping of the long-wavelength components of charge density sometimes impedes convergence. To this end, ABACUS has implemented a modified Kerker preconditioner method from v3.5.0. This method sets a minimum threshold $K_{\rm min}$, safeguarding a minimum proportion of the long-wavelength contribution by setting the prefactor to $\max\{\alpha_{\rho}\frac{q^{2}}{q^{2}+q^{2}_{0}},K_{\rm min}\}$ Zhou et al. (2018).
We tested two preconditioner methods over 50 different cases. As depicted in Fig. 4, the modified Kerker preconditioner reduces the number of convergence steps considerably in most instances, as compared to its traditional preconditioner ker (2024). The modified method proves more effective with lower $\alpha_{\rho}$.
III.2.2 Charge Density Mixing
Compared with non-magnetic calculations, magnetic SCF often struggles to converge and sometimes may even converge to incorrect magnetic configurations. For the Broyden/Pulay mixing method, the residual vector $R$ is formulated as a function of the charge $\rho$ and magnetization densities $\bm{m}$ as
$R(\rho,\bm{m})=R(\rho_{\uparrow}+\rho_{\downarrow},\rho_{\uparrow}-\rho_{\downarrow}).$ (16)
ABACUS v3.5 starts to utilize the mixing (Eq. 17) that leverages the improved residual definition shown in Eq. 16. It is demonstrated in Fig. 5 that using Eq.16 for defining residuals leads to notably better results compared to using the direct charge density $R(\rho_{\uparrow},\rho_{\downarrow})$. The $(i+1)$-th input density can be given by
@@ III.2.2 Charge Density Mixing
here $n$ is the number of historical charge densities considered. $p_{j}$ is the coefficient calculated by the Pulay/Broyden method, and $\alpha_{\rho}\frac{q^{2}}{q^{2}+q^{2}_{\rho}}$ actually represent the modified Kerker preconditioner mentioned in Sec.III.2.1. ABACUS considers the preconditioning and mixing parameters for the magnetic density ($q_{\rho},q_{m}$) separately from those for the charge density ($\alpha_{\rho},\alpha_{m}$). By default, ABACUS only applies a preconditioner to the charge density, not the magnetic density ($q_{m}=0$). Users can enable preconditioning for the magnetic density by setting a non-zero value for $q_{m}$ through the mixing_mag_gg0 keyword. Fig. 5 (a) compares the performance of collinear magnetic calculations in practice among different versions of ABACUS.
ABACUS extends the collinear mixing method mentioned in Sec.III.2.2 to non-collinear calculations by defining a more general residual $R$ as a function of four vectors:
@@
where $\bm{m}_{x}$ is the component of magnetic density $\bm{m}=\rho_{\uparrow}-\rho_{\downarrow}$ in the $x$-direction.
If one is not interested in the energies of a given magnetic configuration but wants to determine the ground state by relaxing the magnetic moments’ directions, the standard Broyden mixing algorithm defined by Eqs. 17 and 18 sometimes fails to find the lowest magnetic configuration. ABACUS implements a promising angle mixing method proposed by Ref. Heide and Ono, 2013, where the residual $R$ is defined as a function of two-dimensional vector $\bm{n}$
$R(\bm{n})=R(\rho,|\bm{m}|),$ (19)
where $|\bm{m}^{i}_{in}|-|\bm{m}^{i}_{out}|$ contributes to the residual vector. At $i$-th iteration, and for each real-space point, $\theta^{i}_{out}$ is determined by the angle between $\bm{m}^{i}_{in}$ and $\bm{m}^{i}_{out}$. The next input $\bm{m}^{i+1}_{in}$ can be generated by following rules: i) keep $|\bm{m}^{i}_{out}|$ unchanged, namely $|\bm{m}^{i+1}_{in}|=|\bm{m}^{i}_{out}|$; ii) set the angle between $\bm{m}^{i+1}_{in}$ and $\bm{m}^{i}_{in}$ as $\theta^{i+1}_{in}=\beta\theta^{i}_{out}$; iii) the new input magnetization lies in the plane spanned by $\bm{m}^{i}_{in}$ and $\bm{m}^{i}_{out}$.
Traditional SCF iterations mix the charge density without involving the density matrix. However, the DFT+U method Cococcioni and De Gironcoli (2005); Anisimov et al. (1993); Anisimov, Aryasetiawan, and Lichtenstein (1997); Dudarev et al. (1998); Jiang et al. (2010) relies on the density matrix to construct the operators, leading to the discrepancy between the density matrix and the charge density, especially far from the ground state. This complicates the convergence process. ABACUS adopts a restart mixing method, which enables the density matrix to be mixed. Based on that, an automatic workflow named the U-Ramping method is realized, where ABACUS increases the U parameter step by step.
Owing to the initial charge density being estimated from atomic wavefunctions or other methods, it is not straightforward to make an analogous assumption for the density matrix. At the very beginning, SCF iterations mix only the charge density. Once a predefined threshold for convergence is achieved, the SCF calculations restart and enable the simultaneous mixing of the density matrix and charge density thereafter. This method has been shown to markedly advance the convergence rate of DFT+U simulations (see Fig. 5 (b)). For cases that are hard to converge, the U-Ramping method may prove beneficial Meredig et al. (2010); Dorado et al. (2013). ABACUS begins with an LDA/PBE calculation that excludes the U term, incorporating it incrementally only after surpassing a specified convergence level, and continues to periodically restart SCF computations until convergent with the complete U value.
Notably, this restart mixing method also enhances the convergence efficiency of other approaches that depend on density matrices, such as hybrid functional method (see Sec. V.5), and AI-assisted functional method DeePKS (see Sec. V.6).

### v3

@@ II.1.2 Exchange-Correlation Functionals
$m(\mathbf{r})=\rho^{\uparrow}(\mathbf{r})-\rho^{\downarrow}(\mathbf{r}).$ (6)
In this regard, the exchange-correlation energy $E_{\rm xc}[\rho^{\uparrow},\rho^{\downarrow}]$ exhibits an explicit functional dependence on $\rho^{\uparrow}$ and $\rho^{\downarrow}$. The corresponding exchange-correlation potential for spin channel $\sigma$ (where $\sigma\in\{\uparrow,\downarrow\}$) is derived via the functional derivative
$V_{\text{xc},\sigma}(\mathbf{r})=\frac{\delta E_{\text{xc}}[\rho^{\uparrow},\rho^{\downarrow}]}{\delta\rho_{\sigma}(\mathbf{r})}.$ (7)
Since the invention of KS-DFT, numerous density functional approximations have been developed. The earliest implementation, the local density approximation (LDA) Kohn and Sham (1965), is expressed as a functional of the spin-summed electron density $\rho(\mathbf{r})$. Subsequent refinements introduced the local spin density approximation (LSDA) Von Barth and Hedin (1972); Vosko, Wilk, and Nusair (1980), which incorporates spin density components $\rho^{\uparrow}$ and $\rho^{\downarrow}$ within a broken-symmetry Slater determinant framework, thereby correctly describing the dissociation limit of H2.
Including density gradients $\nabla\rho(\mathbf{r})$ (or spin density gradients $\nabla\rho^{\uparrow}$ and $\nabla\rho^{\downarrow}$) as additional variables led to the development of generalized gradient approximation (GGA) functionals, some of which substantially improved the KS-DFT accuracy compared to LDA/LSDA. For instance, the PBE functional Perdew, Burke, and Ernzerhof (1996) reduces the mean absolute error (MAE) for molecular atomization energies of 20 main-group molecules from 31.4 kcal/mol (LSDA) to 7.9 kcal/mol.
Further advancement was achieved with meta-GGA functionals, which incorporate the kinetic energy density $\tau(\mathbf{r})$. This class of functionals, exemplified by the strongly constrained and appropriately normed (SCAN) functional Sun, Ruzsinszky, and Perdew (2015), offers enhanced accuracy over GGAs. Variants of SCAN, such as rSCAN Bartók and Yates (2019) and R2SCAN Furness et al. (2020), were subsequently developed to improve numerical stability and convergence while largely retaining its accuracy.
One may also include a fraction of exact Hartree-Fock exchange energy into the exchange-correlation term, and the resulting functional is called hybrid functionals, such as the HSE06 Heyd, Scuseria, and Ernzerhof (2003); Ge and Ernzerhof (2006), PBE0 Adamo and Barone (1999a) and B3LYP functionals Stephens et al. (1994).
Analogous to spin-unpolarized systems, approximations for $E_{\rm xc}[\rho^{\uparrow},\rho^{\downarrow}]$ are typically constructed hierarchically. The simplest approach, the LSDA Miehlich et al. (1989), expresses $E_{\rm xc}$ solely as a function of the local spin densities $\rho^{\sigma}(\mathbf{r})$. More refined treatments, such as the generalized gradient approximation (GGA), introduce additional dependence on the gradients of the spin densities $\nabla\rho^{\sigma}(\mathbf{r})$. Further sophistication is achieved in meta-GGA functionals through the incorporation of the kinetic energy density, while hybrid functionals explicitly include fractional exact exchange integrals.
The LDA, GGA, and meta-GGA functionals have already been implemented in ABACUS for both plane-wave and numerical atomic orbital basis sets. Both LDA and GGA functionals are accessible through either native implementations or the Libxc library Lib (2024), whereas the meta-GGA functional is currently restricted to Libxc. Benchmark studies of SCAN, rSCAN Bartók and Yates (2019), and r2SCAN Furness et al. (2020) within ABACUS are detailed in Ref. Liu et al., 2023. Efficient hybrid functional calculations are available for localized basis sets, utilizing combined capabilities of Libxc and the local resolution of identity (LRI) methodology (see Sec. IV.4 for computational details).
II.1.3 Non-Collinear Spin
@@
The non-collinear spin calculation allows electron spins to orient in arbitrary directions rather than being restricted to a single axis. Within the framework of non-collinear spin KS-DFT, the Kohn-Sham electronic wavefunction $i$ can be represented as two-component spinors as
$\Psi_{i}(\mathbf{r})=\begin{pmatrix}\psi_{i}^{\uparrow}(\mathbf{r})\\ \psi_{i}^{\downarrow}(\mathbf{r})\end{pmatrix}$. (8)
The spin-resolved density matrix $\rho^{\sigma\sigma^{\prime}}$ (with $\sigma,\sigma^{\prime}\in\{\uparrow,\downarrow\}$) contains the full spin structure of the system. It can be further decomposed into charge and spin contributions via the Pauli matrices $\bm{\sigma}=(\sigma_{x},\sigma_{y},\sigma_{z})$ and the identity matrix $\bm{I}$ as follows
@@
$\left(\begin{array}[]{cc}\rho^{\uparrow\uparrow}&\rho^{\uparrow\downarrow}\\ \rho^{\downarrow\uparrow}&\rho^{\downarrow\downarrow}\end{array}\right)$ $=(\rho\bm{I}+\bm{\sigma}\cdot\bm{m})$ (9)
$=\frac{1}{2}\left(\begin{array}[]{cc}\rho+m_{z}&m_{x}-im_{y}\\ m_{x}+im_{y}&\rho-m_{z}\end{array}\right).$
@@
where $\bm{I}$ is a $2\times 2$ identity matrix, $m_{x}$, $m_{y}$, and $m_{z}$ refer to magnetic densities along $x$, $y$, and $z$ axis, respectively.
Next, the exchange-correlation potential $V_{\rm xc}$ adopts a $2\times 2$ matrix form acting on two-component spinors
@@
where the off-diagonal elements $V_{\rm xc}^{\uparrow\downarrow}$ and $V_{\rm xc}^{\downarrow\uparrow}$ account for spin-flip interactions arising from spin-orbit coupling (SOC) or other magnetic effects. These terms mediate the mixing between spin-up and spin-down states, enabling the description of non-collinear magnetism.
@@
Using the Pauli matrix as basis, the exchange-correlation potential can be converted into
$v_{\rm xc}(\mathbf{r})$ $=\frac{\delta E_{\rm xc}}{\delta\rho(\mathbf{r})}=\frac{1}{2}\Bigl{[}V_{\rm xc}^{\uparrow\uparrow}(\mathbf{r})+V_{\rm xc}^{\downarrow\downarrow}(\mathbf{r})\Bigr{]},$ (11)
$b^{x}_{\rm xc}(\mathbf{r})$ $=-\frac{\delta E_{\rm xc}}{\delta m_{x}(\mathbf{r})}=\frac{1}{2}\Bigl{[}V_{\rm xc}^{\uparrow\downarrow}(\mathbf{r})+V_{\rm xc}^{\downarrow\uparrow}(\mathbf{r})\Bigr{]},$
$b^{y}_{\rm xc}(\mathbf{r})$ $=-\frac{\delta E_{\rm xc}}{\delta m_{y}(\mathbf{r})}=\frac{i}{2}\left[V_{\rm xc}^{\uparrow\downarrow}(\mathbf{r})-V_{\rm xc}^{\downarrow\uparrow}(\mathbf{r})\right],$
$b^{z}_{\rm xc}(\mathbf{r})$ $=-\frac{\delta E_{\rm xc}}{\delta m_{z}(\mathbf{r})}=\frac{1}{2}\Bigl{[}V_{\rm XC}^{\uparrow\uparrow}(\mathbf{r})-V_{\rm XC}^{\downarrow\downarrow}(\mathbf{r})\Bigr{]}.$
Here $\mathbf{b}_{xc}$ is the effective exchange-correlation magnetic field, defined as the functional derivative of the exchange-correlation energy with respect to the magnetization density. During Hamiltonian construction, the potential must be reconstructed in its spinor form to operate on the two-component Kohn-Sham wavefunctions.
II.2 Self-Consistent Field Method
@@
Figure 3: ABACUS performs self-consistent field (SCF) calculations with two available basis sets: PW and NAOs. The PW basis set employs iterative diagonalization methods like CG and Davidson to solve the Kohn-Sham equations. For NAOs, it uses exact diagonalization or the low-scaling method PEXSI Lin et al. (2009, 2013). ABACUS applies different mixing algorithms for various types of calculations, including non-magnetic, collinear, and non-collinear. Additionally, ABACUS supports the density matrix mixing method for DFT+U, hybrid functionals, and DeePKS calculations with the NAO basis set.
II.2.1 Mixing of Electron Densities
The Kohn-Sham equations are typically solved through the SCF method, an iterative procedure where the input electron density $\rho_{\text{in}}$ is systematically updated based on various mixing methods until achieving convergence. The efficiency and reliability of this procedure critically depend on the choice of electron density mixing algorithm, which serves the dual purpose of accelerating convergence toward the ground state while preventing stagnation in local minima. The simplest approach, linear (or “plain”) mixing, updates the density for non-spin-polarized systems as
$\rho^{i+1}_{in}=\rho^{i}_{in}+\alpha_{\rho}R[\rho^{i}_{in}],$ (12)
where $i$ denotes the iteration index. $\alpha_{\rho}\in[0,1]$ is the mixing parameter, and a residual vector can be defined as
@@ II.2.1 Mixing of Electron Densities
Besides the linear mixing method, ABACUS implements Pulay’s direct inversion in iterative subspace (DIIS) method Pulay (1980, 1982) and Broyden’s quasi-Newton approach Johnson (1988). Fig. 3 illustrates the electron density mixing strategies implemented in ABACUS to guarantee stable SCF convergence.
Magnetic calculations usually exhibit significantly poorer convergence behavior compared to their non-magnetic counterparts, frequently encountering two distinct failure cases, i.e., divergence of SCF or convergence to excited magnetic configurations. These issues are due to the complex coupling between charge and magnetization degrees of freedom when solving the Kohn-Sham equations. To address these challenges, ABACUS implements a mixing scheme that explicitly accounts for spin polarization. The residual vector $R$ is reformulated to simultaneously track both charge and magnetization density variations
$R(\rho,\bm{m})=R(\rho^{\uparrow}+\rho^{\downarrow},\rho^{\uparrow}-\rho^{\downarrow}).$ (14)
where $\rho$ represents the total charge density and $m$ denotes the magnetization vector. The mixing procedure employs a history-dependent update with wavevector-dependent preconditioning
@@
where $n$ determines the history length, $p_{j}$ are Pulay/Broyden coefficients, and the $q$-dependent terms represents a modified Kerker preconditioner Zhou et al. (2018). The damping parameters $q_{\rho}$ and $q_{m}$ control the suppression of long-wavelength oscillations in charge and magnetization densities, respectively.
ABACUS provides several choices to optimize magnetic calculations, beginning with the implementation of separate preconditioning parameters for charge $q_{\rho}$ and magnetization $q_{m}$ densities to address their distinct physical characteristics and convergence behaviors. By default, the code deactivates magnetic density preconditioning (setting $q_{m}$ = 0) to prioritize numerical stability in most cases, while still providing users the flexibility to enable magnetic preconditioning when enhanced convergence for challenging magnetic systems is required. This approach combines robust default settings with customizable options, allowing the code to automatically handle routine cases while permitting users to fine-tune the mixing parameters for complex magnetic configurations. We compared the magnetic convergence for several magnetic materials using two versions of the ABACUS (v3.4 and v3.5). All computational details are available online mag (2024). As illustrated in Fig. 4(a), a marked difference in SCF convergence behavior is observed between the two versions. Our findings demonstrate that ABACUS v3.5 achieves significantly enhanced convergence efficiency when compared to the version 3.4 of ABACUS.
ABACUS generalizes the collinear mixing formalism to non-collinear magnetic calculations through an expanded residual definition that incorporates all magnetization components
@@
where $\bm{m}_{x}$ denotes the $x$-component of the magnetization density $\bm{m}=\rho^{\uparrow}-\rho^{\downarrow}$, with analogous definitions for the $y$ and $z$ components. While this formulation successfully extends the standard Broyden mixing approach (Eqs. 15 and 16) to non-collinear cases, it may fail to converge to the true ground state when magnetic moment directions require optimization. To address this limitation, ABACUS implements an angle mixing scheme based on the work of Ref. Heide and Ono, 2013. This approach reformulates the residual in terms of a reduced parameter space $R(\rho,|\bm{m}|)$. The residual incorporates the magnitude difference $|\bm{m}_{\text{in}}^{i}|-|\bm{m}_{\text{out}}^{i}|$, where $|\bm{m}_{\text{in}}^{i}|$ is the input spin density of $i$-th iteration and $|\bm{m}_{\text{out}}^{i}|$ is the calculated output one. The mixing algorithm proceeds through three key steps at each iteration $i$. First, $|\bm{m}_{\text{in}}^{i+1}|=|\bm{m}_{\text{out}}^{i}|$ is ensured to keep local moment unchanged. Second, for the angle relaxation, the input angle is computed as $\theta_{\text{in}}^{i+1}=\beta\theta_{\text{out}}^{i}$, where $\theta_{\text{out}}^{i}$ measures the angular deviation between $\bm{m}_{\text{in}}^{i}$ and $\bm{m}_{\text{out}}^{i}$, and $\beta$ serves as a relaxation parameter. Third, $\bm{m}_{\text{in}}^{i+1}$ is restricted to the plane defined by $\bm{m}_{\text{in}}^{i}$ and $\bm{m}_{\text{out}}^{i}$ to maintain geometric consistency. This methodology enables efficient exploration of magnetic configuration space while avoiding common convergence pitfalls associated with traditional non-collinear mixing approaches.